what is the value of log6 1/36
Answer by lwsshak3(11628)

what is the value of log6 1/36
..
log6 1/36
What is the logarithm of 1/36 to the base 6?
Change expression into the exponential form
The base raised to the logaritlhm of the number is equal to the number
6^x=1/36=1/6^2=6^-2
x=-2
The logarithm of 1/36 to the base 6=-2